d²y/d²x+2dy/dx+10y=0


1
Expert's answer
2021-05-02T05:34:10-0400

"\frac{d^{2}y}{dx}+2\frac{dy}{dx}+10y=0"


Let "y=e^{mx}"


Then "\frac{dy}{dx}=me^{mx}" and "\frac{d^{2}y}{dx^{2}}=m^{2}e^{mx}"


Therefore we get "(m^{2}+2m+10)e^{mx}=0"


As "e^{mx}\neq 0," we have "(m^{2}+2m+10)=0"

i.e., "m=\frac{-2\pm \sqrt{ 4-40}}{2}"


"=\frac{-2\pm i6}{2}"

"=-1\pm i3"


Therefore the solution becomes "y=e^{-1}(c_{1} Cos 3x+c_{2} Sin 3x)" where "c_{1},c_{2}" are constants.